OpenStudy (anonymous):

10/31=5W/62 solve the proportion step by step

OpenStudy (anonymous):

is the original equation 10/31=5(W/62)?

OpenStudy (anonymous):

\[\frac{ 10 }{ 31 }=5\frac{ W }{ 62 }\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{ 10 }{ 31}=\frac{ 5W }{ 62 }\]

OpenStudy (anonymous):

so how i wrote is the original way..

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so multiply both sides by the LCD, which is 62 the equation becomes: 20=5W divide both sides by 5 and you get W=4

OpenStudy (anonymous):

i see! thnk u so muchh!